how many moles of water are in a snowflakes weighing 5.0x10^-5g? How many of these snowflakes would it take to have 1.0kg of ice?

1
Expert's answer
2021-11-05T01:07:38-0400

Solution:

The molar mass of water (H2O) is 18.015 g/mol

Hence,

(5.0×10-5 g H2O) × (1 mol H2O / 18.015 g H2O) = 2.775×10-6 mol H2O = 2.8×10-6 mol H2O


Convert kg to g:

(1.0 kg) × (1000 g / 1 kg) = 1000 g


Number of snowflakes = (1000 g) / (5.0×10-5 g) = 2.0×107 snowflakes


Answer:

2.8×10-6 moles of water

2.0×107 snowflakes